OpenStudy (hockeychick23):

Consider a set of 7500 scores on a national test whose score is known to be distributed normally with a mean of 510 and a standard deviation of 85. About how many scores greater than 600 would we expect to find? A. 3750 B. 2664 C. 6414 D. 575 E. 1086

OpenStudy (hockeychick23):

@perl i did 6899/7500 since before you said find the probability of choosing a score greater than 600 but I'm not sure what to do next

OpenStudy (perl):

P( X > 600) = P( Z > ? )

OpenStudy (perl):

first lets change the raw score of 600 to a z score

OpenStudy (perl):

the z score tells you how many standard deviations away you are from the mean . so how many standard deviations away from the mean is the score of 600 ?

OpenStudy (hockeychick23):

600-510/x= 7500 90/x=7500 x= 83.3

OpenStudy (hockeychick23):

i think i did something wrong finding x

OpenStudy (perl):

z = (600 - 510) / 85 will give you the z score

OpenStudy (hockeychick23):

oh sorry thanks, ok so 90/85=1.05882

OpenStudy (perl):

so we want P( Z > 1.0588)

OpenStudy (hockeychick23):

i looked it up on the table and got around .8554

OpenStudy (perl):

the table gave you the value for P( Z < 1.0588) , so we need to subtract that from 1 to get P( Z > 1.0588)

OpenStudy (hockeychick23):

ok so i got 1-.8554=.1446

OpenStudy (perl):

correct

OpenStudy (perl):

now multiply that probability by 7500 , to get how many you expect to score more than 600

OpenStudy (hockeychick23):

ok when i did that i got 1084.5 which is close to one of my answers 1086

OpenStudy (hockeychick23):

thanks!(: